OpenStudy (anonymous):

A 108 meter long piece of wire is connected to a 3.4 volt battery. The cross-sectional area of the wire is 2.270×10-6 m2. The wire is made out of a material having a resistivity of 9.32·10-8 Ω · m. What is the magnitude of the current (in Amperes) flowing through the wire?

OpenStudy (anonymous):

Please help!!!!!!!!!!!!

OpenStudy (anonymous):

I know R=p*L/A --> this equals 4.43E-12 Then I=V/R--> 3.4/4.43E-12 this equals 7.67E11 but it keeps telling me its wrong

OpenStudy (anonymous):

Your approach is correct. Your answer is also correct except the order of magnitude. Actually, R = 4.43 ohm

OpenStudy (anonymous):

Thank yoU!!!! why is it 4.43?

OpenStudy (anonymous):

Could you help me with my next one I am about to post as well? I'd really appreciate it! It's due at 5 and ive been working on it all day

OpenStudy (anonymous):

\(R=\rho L/A =9.32·10^{-8}\ Ω · m \times 108\ m / 2.270×10^{-6} \ m^2 = 443 \times 10^{-8 +6} =443 \times 10^{-2} = 4.43 \)